I want to know what I need to do to solve this thanks.
Henry is three times as old as Jane. In 3 years Henry will be x years old. In terms of x, how old will Jane be then?
Jane's Age then = ?

Answer by whisper(27)

Let's break this up.
Henry = 3 x Jane
Henry + 3 = x
What is Jane + 3?
Let's look at Henry + 3 = x
Subtract 3 from both sides of the equation
Henry + 3 - 3 = x - 3
Henry + 0 = x - 3
Henry = x - 3
But Henry also = 3 x Jane
So we can set
x-3 = 3 x Jane
Divide both sides by 3.
(x-3)/3=(3 x Jane)/3
(x-3)/3= Jane
So Jane is (x-3)/3
But we want Jane + 3
So
[(x-3)/3] + 3 = Jane + 3
To simplify the answer find a common denominator:
(x-3)/3 + 9/3
(x-3 + 9)/3
(x+6)/3 This is ok for an answer or you could write it as:
x/3 + 2 Either is acceptable.
Don't hesitate to ask if you have further questions!